ABC bank produced net income of just over sh.23 billion in 2015, and had total assets of sh.1.787 trillion at the end of the year. Compute its return on assets (ROA).
1
Expert's answer
2018-10-29T16:34:09-0400

ROA=(net income)/(total assets)=23/1787=1.29%